OpenStudy (anonymous):

Solve the equation : 2x-5=3(x+1)

OpenStudy (anonymous):

please help

hartnn (hartnn):

same question again ?

OpenStudy (anonymous):

pooja can u please go

OpenStudy (anonymous):

hartnn is with me

hartnn (hartnn):

lol, anyone can come :) its an open internet

OpenStudy (igreen):

^^^^^^^

OpenStudy (anonymous):

@hartnn when i did the ques again i got the ans as x=8

OpenStudy (anonymous):

let me show u how

hartnn (hartnn):

can u show the steps ?

OpenStudy (igreen):

^^^^^^^^^^^^^^^^^^^^^^^^ I got something else.

OpenStudy (anonymous):

2x-5=3(x+1)

OpenStudy (anonymous):

2x-5=3x+3

OpenStudy (anonymous):

please corrcet me where i am wrong

hartnn (hartnn):

correct, continue

OpenStudy (anonymous):

2x-3x=5+3

OpenStudy (anonymous):

i did it differently

OpenStudy (igreen):

THat's correct..

OpenStudy (anonymous):

is it corrrect

OpenStudy (igreen):

*That's Next step..

OpenStudy (anonymous):

hart u tell me

OpenStudy (anonymous):

is it correct

hartnn (hartnn):

it is correct

OpenStudy (igreen):

You're correct..the next step is to combine like terms. 2x and 3x 3 and 8

OpenStudy (anonymous):

next step is

OpenStudy (igreen):

3 and 5*

OpenStudy (anonymous):

2x-3x=8

OpenStudy (anonymous):

am i correct

OpenStudy (anonymous):

next is

hartnn (hartnn):

yes

OpenStudy (anonymous):

-x=8

OpenStudy (igreen):

Correct..

hartnn (hartnn):

correct

OpenStudy (anonymous):

x=-8

OpenStudy (igreen):

Yep.

hartnn (hartnn):

yesh

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

i am sorry

hartnn (hartnn):

iGreen, you type quite fast :)

OpenStudy (anonymous):

i think i messed up a little

hartnn (hartnn):

don't be sorry, its ok :)

OpenStudy (anonymous):

bye

OpenStudy (igreen):

Up to 100WPM ;)

hartnn (hartnn):

lol

OpenStudy (anonymous):

LOOK before u leap.lol

OpenStudy (anonymous):

do you still need help?